ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

101591991 MS Classification: 05, 52 Gawrilow, Ewgenij, Joswig, Michael, Technische Universität Berlin, Germany polymake: a Software Environment for the Investigation of Polytopes Polymake is a free software package for the algorithmic treatment of polytopes. It offers a common interface to many other packages developed by other research groups. For polymake, a polytope can be specified in many difrm31. txtferent ways. Then a great variety of questions about (data and properties of) the polytope can be asked. Polymake will provide and execute a suitable sequence of algorithms in order to supply the answer. Polymake is intended for researchers who are interested in polyhedral geometry and linear optimization. At the same time polymake is used for students’ education.

See http://www.math.tu-berlin.de/diskregeom/polymake/doc/ for de-

tails.